<html>
<head>
<meta http-equiv="Content-Type" content="text/html; charset=iso-8859-1">
<style type="text/css" style="display:none;"> P {margin-top:0;margin-bottom:0;} </style>
</head>
<body dir="ltr">
<div style="text-align: left; text-indent: 0px; margin-top: 0.4em; margin-bottom: 0.8em; font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Dear Collaborators,</div>
<div style="text-align: left; text-indent: 0px; margin-top: 0.4em; margin-bottom: 0.8em; font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Yesterday was RF recovery day for the accelerator and the hall was in controlled access. The survey was scheduled to take place at 10:00, but issues with PSS in hall C and a dead wand from RADCON delayed it some. The hall was in controlled access by around
11:15, and we locked up and returned to beam permit around 2:25. During this time, teams mostly from Moller made use of the access. The target group did not make an access as they were occupied elsewhere, and anyways indicated there was little they could do
without major work. For now, however, it is okay to keep running beam. From SBS side, the GEM team were the only ones to make an access to add additional gas lines to some modules. They would later continue their tests when beam came back (<span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWA9a6cc2f1-8b7e-595f-8417-e65d9b0ff76e" href="https://logbooks.jlab.org/entry/4447500">https://logbooks.jlab.org/entry/4447500</a></span>).
Beam returned around 5:00, but early harp scans looked off, and it was realized there were issues with the raster (<span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWA3968cf7e-f698-0360-2f18-aae83f39bb39" href="https://logbooks.jlab.org/entry/4447030">https://logbooks.jlab.org/entry/4447030</a></span>).
On shift, Silviu still noticed issues with the raster (<span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWA91df8694-15b9-96b8-26db-1fe9bddbf0c6" href="https://logbooks.jlab.org/entry/4447092">https://logbooks.jlab.org/entry/4447092</a></span>),
MCC indicated this morning they would look into it further. Finally around 8:30 cross centering could be completed and by 8:40 we were back to taking production data. ECal team implemented increases to both the ECal (up to 75%) and HCal (up to 50%) thresholds,
allowing us to push the beam current up to 20 uA with a live time of ~93% (<span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWAfed7a6db-476e-0de8-ea5f-bf6a2d699575" href="https://logbooks.jlab.org/entry/4447530">https://logbooks.jlab.org/entry/4447530</a></span>).
Production ran through the night with these conditions. The Compton test scheduled for today was canceled due to issues with the Compton laser (<span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWA93697f0b-20b4-885b-5a02-fefb9a2c8c20" href="https://logbooks.jlab.org/entry/4446889">https://logbooks.jlab.org/entry/4446889</a></span>),
so we will continue production through the day. The current run plan is to continue taking production data at 20 uA on LH2.</div>
<div style="text-align: left; text-indent: 0px; margin-top: 0.4em; margin-bottom: 0.8em; font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
There was an official announcement this morning at the MCC meeting: Unfortunately, there will be no extension for Hall A. Beam will be taken away at 7:00 a.m. this coming Monday, the 25th.</div>
<div style="text-align: left; text-indent: 0px; margin-top: 0.4em; margin-bottom: 0.8em; font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Since yesterday's Day to today's owl:<br>
* Total Accumulated Charge at Kin3 so far: 88.30 C (Goal: 100 C)<br>
* Total Accumulated Charge at Kin3 since yesterday : 0.69 C</div>
<div style="text-align: left; text-indent: 0px; margin-top: 0.4em; margin-bottom: 0.8em; font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Updates from MCC:<br>
* The Energy Secretary tour is today. Hall C is now in restricted access, to be toured around 10:40. Hall B will be surveyed around 11:30 for the tour at 14:35. MCC tour will take place at 14:10. There may be a short interruption to be beam around 14:10 only<br>
* There were issues with 5th pass vertical separation during recovery yesterday, MCC may take away beam today for a couple hours today (ideally while Hall C and B are offline)<br>
* Compton test was canceled</div>
<div style="text-align: left; text-indent: 0px; margin-top: 0.4em; margin-bottom: 0.8em; font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Current Issues and Hall activities:<br>
* ECal/HCal threshold adjustments<br>
* Ongoing GEM gas tests<br>
* Target leak/pump failure, still holding steady around 2e-6 level</div>
<div style="text-align: left; text-indent: 0px; margin-top: 0.4em; margin-bottom: 0.8em; font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Shift Summary:<br>
* Yesterday's Day: <span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWA6bd3755b-6844-b717-31de-15f8dfa359f7" href="https://logbooks.jlab.org/entry/4446687">https://logbooks.jlab.org/entry/4446687</a></span><br>
* Yesterday's Swing: <span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWA0663b8c7-b5ba-160e-63b1-249582372a82" href="https://logbooks.jlab.org/entry/4446968">https://logbooks.jlab.org/entry/4446968</a></span><br>
* Today's Owl: <span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWAb4868df5-faa9-14b3-4525-ed4293aaa299" href="https://logbooks.jlab.org/entry/4447251">https://logbooks.jlab.org/entry/4447251</a></span></div>
<div style="text-align: left; text-indent: 0px; margin-top: 0.4em; margin-bottom: 0.8em; font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
RC Meeting:<br>
* We will have our usual RC meeting today at 1 pm in the conference room in the second floor of the CH and on Zoom<br>
* <span style="color: rgb(0, 0, 238);"><a style="color: rgb(0, 0, 238);" class="OWAAutoLink" id="OWAe1d681f9-db04-23eb-5dc6-7582be136341" href="https://urldefense.proofpoint.com/v2/url?u=https-3A__jlab-2Dorg.zoomgov.com_j_1616220064&d=DwMFAw&c=CJqEzB1piLOyyvZjb8YUQw&r=cE_5H9I241Ivk1xpqB8Fq0UiLTkf4zMIQ6sedLYb2m8&m=PU4Y-xiThmB-pMmbkiaOcLMosjiQ4dPcUt8fsj_m91KZ4M7DBqjfnEKl2651Gycv&s=LlAHGtQJsNgjmNdonvFnfEkKl04u20H-pKIeyPai5aY&e=">https://jlab-org.zoomgov.com/j/1616220064</a></span><br>
* Meeting ID: 161 622 0064<br>
* Passcode: 314159</div>
<div style="font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Best,</div>
<div style="font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Jacob McMurtry</div>
<div style="font-family: Aptos, Aptos_EmbeddedFont, Aptos_MSFontService, Calibri, Helvetica, sans-serif; font-size: 12pt; color: rgb(0, 0, 0);" class="elementToProof">
Hall A RC</div>
</body>
</html>